OpenStudy (anonymous):

find the value of the annuity. a1=$7000 I=0.10 n=5

OpenStudy (anonymous):

@jim_thompson5910 I followed the steps but when it comes to finding the value I dont get it. This is where I have gotten so far. \[7000\frac{ (1 + 0.10)^5 - 1 }{ 0.10 }\]

jimthompson5910 (jim_thompson5910):

assuming you're using the correct formula

OpenStudy (anonymous):

\[this is the formula I was told \to use: Sn = a1( \frac{ (1 + I)^n -1 }{ I })\]

jimthompson5910 (jim_thompson5910):

\[\large 7000\frac{ (1 + 0.10)^5 - 1 }{ 0.10 }\] \[\large 7000\frac{ (1.10)^5 - 1 }{ 0.10 }\] \[\large 7000\frac{1.61051 - 1 }{ 0.10 }\] \[\large 7000\frac{0.61051 }{ 0.10 }\] \[\large 7000*6.1051\] \[\large 42,735.70\]

OpenStudy (anonymous):

oh and round to the nearest cent

jimthompson5910 (jim_thompson5910):

I think that's right. Does it say anything about when it's compounded?

OpenStudy (anonymous):

No it doesn't.

jimthompson5910 (jim_thompson5910):

alright

jimthompson5910 (jim_thompson5910):

oh and does it say anything about monthly payments?

OpenStudy (anonymous):

No it does not. The answer you gave me is correct. I just checked it :)

jimthompson5910 (jim_thompson5910):

oh, just checking not very good with finance problems lol

jimthompson5910 (jim_thompson5910):

but glad it works out

OpenStudy (anonymous):

You've helped me so much though!!

jimthompson5910 (jim_thompson5910):

glad to be of help